(a)   Ministry of Higher Education (MOHE) put a guideline that maximum of 30% students are allowed to be in campus to control the spread of Covid-19. In average, 13 students will come to Statistics class with standard deviation of 3 students. A sample of 20 students are chosen. Find probability that the average number of students come to Statistics class is less than 15?


1
Expert's answer
2021-08-08T17:21:11-0400

Solution:

"P( 
X

 <15)=P(Z< \dfrac{15-13}{\dfrac{3}{\sqrt{20}}}
 )=P(Z<2.98)=0.9986"